Compute the prices and yields of the following bonds: Issuer Volkswagen JP Morgan Renault Settiement Today Today Today Coupon rate 2.75% 2.25% 0% Frequency coupon payment Term to Maturity Face value Yield to Maturity Price Current Yield Annual Semi-annual Zero-coupon 2 years20 years 5 years $1,000 3.4% 100 4.2% 100 5.0%
Answer #1

Assumption: The coupon rate given is per annum.

1) Volkswagen

Calculation of price of bond:

Price of bond =  c × F × {1 − (1 + r)-t ] / r } + [ F / (1 + r)t]

= 0.0275 * { 1- (1+ 0.05)-2 / 0.05 } + [ 100 / (1+0.05)2 ]

= Euro 95.82

Calculation of current yield:

Current Yield = Annual coupons / Bond price * 100

= 2.75 /95.82 * 100

= 2.869 %

2) JP Morgan :

Calculation of price of bond:

Price of bond =  c × F × {1 − (1 + r)-t ] / r } + [ F / (1 + r)t]

= 0.0225 * { 1- (1+ 0.034)-20 / 0.034 } + [ 100 / (1+0.0034)20 ]

= $ 834.10

Calculation of current yield:

Current Yield = Annual coupons / Bond price * 100

= 45 / 834.10 * 100

= 5.40%

3) Renault :

  • Calculation of price of bond:

Zero Coupon Bond Value = (1r) F-face ualue of bond r-rate or yield t-time t o maturity

Therefore,

Price = 100 / (1+0.042)5

   = 100/1.22839

= 81.406 Euro

  • Calculation of current yield:

Current yield = Annual coupons / Bond price * 100

= 0 /81.40*100

= 0